Use the compound interest formula A=P(1+r)^t and the given information to solve for r.
A=9,000,000
P=60,000
t=40
Answer #1
Just plug in the values

60000(1+r)^40 = 9000000
(1+r)^40 = 150
take 40th root of both sides
1+r = 150^(1/40) = 1.13345
r = .13345 or 13.345%
answered by: mTaq14